OpenStudy (anonymous):

Solve: b^2+2b+4=0 Please show work! Thank you! (:

OpenStudy (anonymous):

Solve it using quadratic equations :)

OpenStudy (anonymous):

or even complete the square

OpenStudy (anonymous):

you will get complex zeros, is that ok? i mean they will not be real

OpenStudy (anonymous):

It would be a complex Nos..!!

OpenStudy (anonymous):

\[b^2+2b+4=0\] \[b^2+2b=-4\] \[(b+1)^2=-4+1=-3\] \[b+1=\pm\sqrt{-3}=\pm\sqrt{3}i\] \[b=-1\pm\sqrt{3}i\]

OpenStudy (anonymous):

\[(-2\pm \sqrt{-12})/2\]

OpenStudy (anonymous):

Thank you both! :D